# Quantum criticality and Kibble-Zurek scaling in the Aubry-André-Stark model

### v1

@@ I Introduction
In recent years, there has been increasing interest in the studies of Anderson localization Anderson (1958); Abrahams et al. (1979); Lee and Ramakrishnan (1985) and localization transitions in quasiperiodic systems Harper (1955); Aubry and André (1980); Lellouch and Sanchez-Palencia (2014); Devakul and Huse (2017a); Roy et al. (2022a); Agrawal et al. (2020); Goblot et al. (2020a); Roy et al. (2022b); Agrawal et al. (2022). Compared to random systems with quenched disorders, quasiperiodic systems exhibit unique localization properties that have been explored both theoretically Goblot et al. (2020a); Agrawal et al. (2020); Roy et al. (2022b); Agrawal et al. (2022) and experimentally Crespi et al. (2013); Bordia et al. (2016); Roati et al. (2008). The one-dimensional Aubry-André (AA) model Aubry and André (1980) serves as an important example in this regard, where a localization transition occurs when the strength of the quasiperiodic potential exceeds the critical point determined by the self-duality Biddle and Das Sarma (2010); Liu et al. (2020); Wang et al. (2023a). Furthermore, various extensions of the AA model have been proposed to investigate the mobility edges Soukoulis and Economou (1982); Ganeshan et al. (2015), topological phases Zhang et al. (2018, 2020); Yoshii et al. (2021); Zhang et al. (2021a); Tang et al. (2022); Wang et al. (2022); Nakajima et al. (2021); Wu et al. (2022); Huang et al. (2024), many-body localization Schreiber et al. (2015); Lukin et al. (2019); Wang et al. (2023b), and critical phenomena Goblot et al. (2020b); Lv et al. (2022a, b); Aramthottil et al. (2021); Devakul and Huse (2017b); Bu et al. (2022, 2023). Remarkably, the quantum criticality and scaling functions with new critical exponents for the localization transition in the disordered AA model have been unearthed in Ref. Bu et al. (2022), where the random disorder contributes an independent relevant direction near the AA critical point. According to the renormalization-group theory Belitz and Vojta (2005); Fisher (1974); Gosselin et al. (2001); Zhong (2006), critical exponents in the scaling functions of physical observables around the critical point Cherroret et al. (2014); Lemarié et al. (2009); Slevin and Ohtsuki (2009); You and Dong (2011); Su and Wang (2018); Wang et al. (2024); Slevin and Ohtsuki (1997); Asada et al. (2002); Rams and Damski (2011) characterize the universal features of continuous quantum (classical) phase transitions Sachdev (2011); Osterloh et al. (2002); Heyl (2018); Carollo et al. (2020); Vojta (2003). Thus, determining critical exponents is crucial in understanding critical phenomena and phase transitions, including the localization transition.
@@
On the other hand, the Kibble-Zurek mechanism Kibble (1976); Zurek (1985); Kibble (1980); Zurek (1996) provides a powerful framework to investigate critical dynamics of phase transitions ranging from cosmology to condensed matter systems Laguna and Zurek (1997); Zurek et al. (2005); Polkovnikov (2005); Dziarmaga (2010); Polkovnikov et al. (2011); Qiu et al. (2020a); Anquez et al. (2016); Gao et al. (2017); Huang et al. (2014); Qiu et al. (2020b). Based on this framework, driven dynamics across a critical point can be described by the universal Kibble-Zurek scaling (KZS) and the critical exponents associated with the phase transition can be extracted. Recently, more and more attention has been paid to the driven dynamics in localization transitions Morales-Molina et al. (2014); Serbyn et al. (2014); Sinha et al. (2019); Sun et al. (2022); Tong et al. (2021); Decker et al. (2020); Bu et al. (2023); Zhai et al. (2022). In particular, the KZS has been generalized to characterize the driven dynamics in the disordered AA model Bu et al. (2022), which can include two scaling variables when both the random and quasiperiodic potentials are relevant directions Bu et al. (2023). Notably, random and quasiperiodic disorders are not the only route to induce localization transitions. For instance, localization can also manifest in systems featuring a linear potential, known as the Wannier-Stark localization in the noninteracting case Wannier (1960); Emin and Hart (1987); Zhang et al. (2021b); Taylor et al. (2020); Lang et al. (2022); Schmidt et al. (2018); Guo et al. (2021a). In the presence of interactions, the Stark many-body localization has been revealed Schulz et al. (2019); Van Nieuwenburg et al. (2019); Guo et al. (2021b); Wang et al. (2021); Morong et al. (2021); Wei et al. (2022); Van Nieuwenburg et al. (2019). It has been shown that the Stark field can induce to diffusive dynamics under the interplay between Anderson and Stark localizations in two-dimensional random lattices Kolovsky (2008a). Moreover, the weak-field sensing with super-Heisenberg precision based on the Stark localization has been proposed He et al. (2023). However, the critical properties and related KZS near localization transitions in the presence of the Stark and quasiperiodic fields remain unexplored.
@@ III.1 Pure Aubry-André and Stark criticalities
The numerical results before and after rescaling $\mathcal{I}$ and $\delta$ as $\mathcal{I}L^{s_{\delta}/v_{\delta}}$ and $\delta L^{1/L^{v_{\delta}}}$ are shown in Figs. 2(b1) and  2(b2), respectively. The best fitting for all the curves is obtained when $s_{\delta}=0.333$. By combing Eqs. (8,9,10) with $\rho=-z_{\delta}v_{\delta}$, we obtain the scaling function of the energy gap $\Delta E$ as
@@
The numerical results of $\xi$ versus the Stark field strength $\varepsilon$ for various $L$ are shown in Fig. 3(a1). One can see that an initial flat region of $\varepsilon$ exists due to the delocalization nature of the wave function for finite $L$, which tends to be smaller (vanishing) as $L$ is increased. The localization length becomes independent on the system size beyond a small threshold, which indicates the presence of the Stark localization. By rescaling $\xi$ and $\varepsilon$ as $\xi L^{-1}$ and $\varepsilon L^{1/\nu_{\varepsilon}}$ in Fig. 3(a2), we find the best collapse for $\nu_{\varepsilon}=0.33$. Similarly, the IPR $\mathcal{I}$ in the pure Stark model satisfies the scaling form
@@ IV KZS of driven dynamics
where $g$ can be $\delta$ or $\varepsilon$ depending on the model considered, and $g_{0}>0$ represents the initial distance from the critical point at $t=0$. According to the KZS, when $|g|>R^{1/\nu r}$ with the scaling exponent $r=z+1/\nu$, the system has enough time to adjust the change of the Hamiltonian to preserve adiabatic; while when $|g|<R^{1/\nu r}$, the change rate of the system itself is less than that of the external parameter, which implies the system entering the impulse region. Below, we choose the system size $L=1597$, which is sufficiently large to ignore the finite-size effect in real-time simulations. We set the initial state as the ground state of the system with fixed $\phi=0.2$ and $g_{0}=1.5$.
@@
where $r_{\delta}=z_{\delta}+1/\nu_{\delta}$ is denoted for the pure AA model. The numerical result of $\xi$ versus $\delta$ for different driving rates $R$ is shown in Fig. 8(a1). We find that when $\delta>R^{1/\nu_{\delta}r_{\delta}}$, $\xi$ is independent of $R$, such that the curves under different $R$ coincide. This implies the adiabatic evolution of the system at this time. When $\delta<R^{1/\nu_{\delta}r_{\delta}}$, the system enters the impulse region and the curves under different $R$ are separated from each other. By rescaling $\xi$ and $\delta$ as $\xi R^{1/r_{\delta}}$ and $\delta R^{-1/\nu_{\delta}r_{\delta}}$ in Fig. 8(a2), we find that the curves of different $R$ collapse well near the critical point, which confirms the KZS form given by Eq. (25). Similarly, the IPR $\mathcal{I}$ satisfies the KZS form
${\mathcal{I}}(\delta,R)=R^{s_{\delta}/r_{\delta}\nu_{\delta}}f_{15}(\varepsilon R^{-1/r_{\delta}\nu_{\delta}}).$ (26)
As shown in Fig. 8(b1), when $\delta>R^{1/\nu_{\delta}r_{\delta}}$ ($\delta<R^{1/\nu_{\delta}r_{\delta}}$ near the critical point) for the adiabatic (impulse) region, the curves of $\mathcal{I}$ for different driving rates $R$ coincide (are separated from each other). By rescaling $\mathcal{I}$ and $\delta$ as $\mathcal{I}R^{-s_{\delta}/\nu_{\delta}r_{\delta}}$ and $\delta R^{-1/\nu_{\delta}r_{\delta}}$ in Fig. 8(b2), we find that the curves collapse according to Eq. (26).
@@
respectively. The numerical results of $\xi$ and $\mathcal{I}$ versus $\varepsilon$ for different driving rate $R$ are shown in Fig. 9(a1) and 9(b1), respectively. The initial evolutions of $\xi$ and $\mathcal{I}$ are independent of $R$ when $\varepsilon>R^{1/\nu_{\varepsilon}r_{\varepsilon}}$. Near the critical point with $\varepsilon<R^{1/\nu_{\varepsilon}r_{\varepsilon}}$, the system enters the impulse region, and the curves of $\xi$ and $\mathcal{I}$ for different driving rate $R$ are separated from each other. By using the collapse plot of $\xi R^{1/r_{\varepsilon}}$ ($\mathcal{I}R^{-s_{\varepsilon}/\nu_{\varepsilon}r_{\varepsilon}}$) and $\varepsilon R^{-1/\nu_{\varepsilon}r_{\varepsilon}}$ ($\varepsilon R^{-1/\nu_{\varepsilon}r_{\varepsilon}}$) in Fig. 9(a2) [Fig. 9(b2)], we confirm the KZS of $\xi$ ($\mathcal{I}$) in Eq. (27) [Eq. (28)] for the pure Stark model with the critical exponents $\nu_{\varepsilon}=0.33$ and $z_{\varepsilon}=2$.
@@
The numerical results of $\xi$ and $\mathcal{I}$ versus $\varepsilon$ for various driving rates $R$ are shown in Figs. 10(a1) and 10(b1), respectively. One can observe that the curves under different $R$ separated from each near the critical point. After rescaling the curves with the critical exponents in AAS model, we find the rescaled curves collapses with each other near the critical point, as shown in Figs. 10(a2) and 10(b2). For $\delta\neq 0$, we can fix the value of $\delta R^{-1/\nu_{\delta}r}$ to verify the KZS forms of $\xi$ and $\mathcal{I}$, as shown in Fig. 11. Here we set $\delta R^{-1/\nu_{\delta}r}=0.3$ and numerically calculate the time evolutions of $\xi$ and $\mathcal{I}$ for various driving rate $R$, with results shown in Figs. 11(a1) and 11(b1). After rescaling of the physical quantities according to Eq. (31) and Eq. (32), one can see that the curves under different $R$ collapse together near the critical point in Figs. 11(a2) and 11(b2). This demonstrates the KZS near the AAS critical point. We also numerically confirm the KZS for the AAS model with fixed $\delta R^{-1/\nu_{\delta}r}=-0.3$.
Figure 11: KZS of driven dynamics in the AAS model with fixed $\delta R^{-1/\nu_{\delta}r}=0.3$. Curves of $\xi$ versus $\varepsilon$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\mathcal{I}$ versus $\varepsilon$ before (b1) and after (b2) rescaling for different $R$. The black arrows denote the quench direction.

### v2

@@ I Introduction
In recent years, there has been increasing interest in the studies of Anderson localization Anderson (1958); Abrahams et al. (1979); Lee and Ramakrishnan (1985) and localization transitions in quasiperiodic systems Harper (1955); Aubry and André (1980); Lellouch and Sanchez-Palencia (2014); Devakul and Huse (2017a); Roy et al. (2022a); Agrawal et al. (2020); Goblot et al. (2020a); Roy et al. (2022b); Agrawal et al. (2022). Compared to random systems with quenched disorders, quasiperiodic systems exhibit unique localization properties that have been explored both theoretically Goblot et al. (2020a); Agrawal et al. (2020); Roy et al. (2022b); Agrawal et al. (2022) and experimentally Crespi et al. (2013); Bordia et al. (2016); Roati et al. (2008). The one-dimensional Aubry-André (AA) model Aubry and André (1980) serves as an important example in this regard, where a localization transition occurs when the strength of the quasiperiodic potential exceeds the critical point determined by the self-duality Biddle and Das Sarma (2010); Liu et al. (2020); Wang et al. (2023a). Furthermore, various extensions of the AA model have been proposed to investigate the mobility edges Soukoulis and Economou (1982); Ganeshan et al. (2015), topological phases Zhang et al. (2018, 2020); Yoshii et al. (2021); Zhang et al. (2021a); Tang et al. (2022); Wang et al. (2022); Nakajima et al. (2021); Wu et al. (2022); Huang et al. (2024), many-body localization Schreiber et al. (2015); Lukin et al. (2019); Wang et al. (2023b), and critical phenomena Goblot et al. (2020b); Lv et al. (2022a, b); Aramthottil et al. (2021); Devakul and Huse (2017b); Bu et al. (2022, 2023); Roósz (2024). Remarkably, the quantum criticality and scaling functions with new critical exponents for the localization transition in the disordered AA model have been unearthed in Ref. Bu et al. (2022), where the random disorder contributes an independent relevant direction near the AA critical point. According to the renormalization-group theory Belitz and Vojta (2005); Fisher (1974); Gosselin et al. (2001); Zhong (2006), critical exponents in the scaling functions of physical observables around the critical point Cherroret et al. (2014); Lemarié et al. (2009); Slevin and Ohtsuki (2009); You and Dong (2011); Su and Wang (2018); Wang et al. (2024); Slevin and Ohtsuki (1997); Asada et al. (2002); Rams and Damski (2011) characterize the universal features of continuous quantum (classical) phase transitions Sachdev (2011); Osterloh et al. (2002); Heyl (2018); Carollo et al. (2020); Vojta (2003). Thus, determining critical exponents is crucial in understanding critical phenomena and phase transitions, including the localization transition.
@@
On the other hand, the Kibble-Zurek mechanism Kibble (1976); Zurek (1985); Kibble (1980); Zurek (1996) provides a powerful framework to investigate critical dynamics of phase transitions ranging from cosmology to condensed matter systems Laguna and Zurek (1997); Zurek et al. (2005); Polkovnikov (2005); Dziarmaga (2010); Polkovnikov et al. (2011); del Campo and Zurek (2014); Das et al. (2012); Bando et al. (2020); Zeng et al. (2023); Qiu et al. (2020a); Anquez et al. (2016); Gong et al. (2016); Gao et al. (2017); Huang et al. (2014); Qiu et al. (2020b). Based on this framework, driven dynamics across a critical point can be described by the universal Kibble-Zurek scaling (KZS) and the critical exponents associated with the phase transition can be extracted. Recently, more and more attention has been paid to the dynamics in disorder-driven transitions Liu et al. (2015); Roósz et al. (2014); Morales-Molina et al. (2014); Serbyn et al. (2014); Sinha et al. (2019); Sun et al. (2022); Tong et al. (2021); Decker et al. (2020); Bu et al. (2023); Zhai et al. (2022). In particular, the KZS has been generalized to characterize the driven dynamics in the disordered AA model Bu et al. (2022), which can include two scaling variables when both the random and quasiperiodic potentials are relevant directions Bu et al. (2023). Notably, random and quasiperiodic disorders are not the only route to induce localization transitions. For instance, localization can also manifest in systems featuring a linear potential, known as the Wannier-Stark localization in the noninteracting case Wannier (1960); Emin and Hart (1987); Zhang et al. (2021b); Taylor et al. (2020); Lang et al. (2022); Schmidt et al. (2018); Guo et al. (2021a). In the presence of interactions, the Stark many-body localization has been revealed Schulz et al. (2019); Van Nieuwenburg et al. (2019); Guo et al. (2021b); Wang et al. (2021); Morong et al. (2021); Wei et al. (2022); Van Nieuwenburg et al. (2019). It has been shown that the Stark field can induce to diffusive dynamics under the interplay between Anderson and Stark localizations in two-dimensional random lattices Kolovsky (2008a). Moreover, the weak-field sensing with super-Heisenberg precision based on the Stark localization has been proposed He et al. (2023). However, the critical properties and related KZS near localization transitions in the presence of the Stark and quasiperiodic fields remain unexplored.
@@ III.1 Pure Aubry-André and Stark criticalities
The numerical results before and after rescaling $\mathcal{I}$ and $\delta$ as $\mathcal{I}L^{s_{\delta}/v_{\delta}}$ and $\delta L^{1/v_{\delta}}$ are shown in Figs. 2(b1) and  2(b2), respectively. The best fitting for all the curves is obtained when $s_{\delta}=0.333$. By combing Eqs. (8,9,10) with $\rho=-z_{\delta}v_{\delta}$, we obtain the scaling function of the energy gap $\Delta E$ as
@@
The numerical results of $\xi$ versus the Stark field strength $\varepsilon$ for various $L$ are shown in Fig. 3(a1). One can see that an initial flat region of $\varepsilon$ exists due to the delocalization nature of the wave function for finite $L$, which tends to be smaller (vanishing) as $\varepsilon$ is increased. The localization length becomes independent on the system size beyond a small threshold, which indicates the presence of the Stark localization. By rescaling $\xi$ and $\varepsilon$ as $\xi L^{-1}$ and $\varepsilon L^{1/\nu_{\varepsilon}}$ in Fig. 3(a2), we find the best collapse for $\nu_{\varepsilon}=0.33$. Similarly, the IPR $\mathcal{I}$ in the pure Stark model satisfies the scaling form
@@ IV KZS of driven dynamics
where $g$ can be $\delta$ or $\varepsilon$ depending on the model considered, and $g_{0}>0$ represents the initial distance from the critical point at $t=0$. According to the KZS, when $|g|>R^{1/r\nu}$ with the scaling exponent $r=z+1/\nu$, the system has enough time to adjust the change of the Hamiltonian to preserve adiabatic; while when $|g|<R^{1/r\nu}$, the change rate of the system itself is less than that of the external parameter, which implies the system entering the impulse region. Below, we choose the system size $L=1597$, which is sufficiently large to ignore the finite-size effect in real-time simulations. We set the initial state as the ground state of the system with fixed $\phi=0.2$ and $g_{0}=1.5$.
@@
where $r_{\delta}=z_{\delta}+1/\nu_{\delta}$ is denoted for the pure AA model. The numerical result of $\xi$ versus $\delta$ for different driving rates $R$ is shown in Fig. 8(a1). We find that when $\delta>R^{1/r_{\delta}\nu_{\delta}}$, $\xi$ is independent of $R$, such that the curves under different $R$ coincide. This implies the adiabatic evolution of the system at this time. When $\delta<R^{1/r_{\delta}\nu_{\delta}}$, the system enters the impulse region and the curves under different $R$ are separated from each other. By rescaling $\xi$ and $\delta$ as $\xi R^{1/r_{\delta}}$ and $\delta R^{-1/r_{\delta}\nu_{\delta}}$ in Fig. 8(a2), we find that the curves of different $R$ collapse well near the critical point, which confirms the KZS form given by Eq. (25). Similarly, the IPR $\mathcal{I}$ satisfies the KZS form
${\mathcal{I}}(\delta,R)=R^{s_{\delta}/r_{\delta}\nu_{\delta}}f_{15}(\delta R^{-1/r_{\delta}\nu_{\delta}}).$ (26)
As shown in Fig. 8(b1), when $\delta>R^{1/r_{\delta}\nu_{\delta}}$ ($\delta<R^{1/r_{\delta}\nu_{\delta}}$ near the critical point) for the adiabatic (impulse) region, the curves of $\mathcal{I}$ for different driving rates $R$ coincide (are separated from each other). By rescaling $\mathcal{I}$ and $\delta$ as $\mathcal{I}R^{-s_{\delta}/r_{\delta}\nu_{\delta}}$ and $\delta R^{-1/r_{\delta}\nu_{\delta}}$ in Fig. 8(b2), we find that the curves collapse according to Eq. (26).
@@
respectively. The numerical results of $\xi$ and $\mathcal{I}$ versus $\varepsilon$ for different driving rate $R$ are shown in Fig. 9(a1) and 9(b1), respectively. The initial evolutions of $\xi$ and $\mathcal{I}$ are independent of $R$ when $\varepsilon>R^{1/r_{\varepsilon}\nu_{\varepsilon}}$. Near the critical point with $\varepsilon<R^{1/r_{\varepsilon}\nu_{\varepsilon}}$, the system enters the impulse region, and the curves of $\xi$ and $\mathcal{I}$ for different driving rate $R$ are separated from each other. By using the collapse plot of $\xi R^{1/r_{\varepsilon}}$ ($\mathcal{I}R^{-s_{\varepsilon}/r_{\varepsilon}\nu_{\varepsilon}}$) and $\varepsilon R^{-1/r_{\varepsilon}\nu_{\varepsilon}}$ ($\varepsilon R^{-1/r_{\varepsilon}\nu_{\varepsilon}}$) in Fig. 9(a2) [Fig. 9(b2)], we confirm the KZS of $\xi$ ($\mathcal{I}$) in Eq. (27) [Eq. (28)] for the pure Stark model with the critical exponents $\nu_{\varepsilon}=0.33$ and $z_{\varepsilon}=2$.
@@
Figure 11: KZS of driven dynamics in the AAS model with fixed $\delta R^{-1/r\nu_{\delta}}=0.3$. Curves of $\xi$ versus $\varepsilon$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\mathcal{I}$ versus $\varepsilon$ before (b1) and after (b2) rescaling for different $R$. The black arrows denote the quench direction.
@@
The numerical results of $\xi$ and $\mathcal{I}$ versus $\varepsilon$ for various driving rates $R$ are shown in Figs. 10(a1) and 10(b1), respectively. One can observe that the curves under different $R$ separated from each near the critical point. After rescaling the curves with the critical exponents in AAS model, we find the rescaled curves collapses with each other near the critical point, as shown in Figs. 10(a2) and 10(b2). For $\delta\neq 0$, we can fix the value of $\delta R^{-1/r\nu_{\delta}}$ to verify the KZS forms of $\xi$ and $\mathcal{I}$, as shown in Fig. 11. Here we set $\delta R^{-1/r\nu_{\delta}}=0.3$ and numerically calculate the time evolutions of $\xi$ and $\mathcal{I}$ for various driving rate $R$, with results shown in Figs. 11(a1) and 11(b1). After rescaling of the physical quantities according to Eq. (31) and Eq. (32), one can see that the curves under different $R$ collapse together near the critical point in Figs. 11(a2) and 11(b2). This demonstrates the KZS near the AAS critical point. We also numerically confirm the KZS for the AAS model with fixed $\delta R^{-1/r\nu_{\delta}}=-0.3$.